A company has decided to produce a new cereal called ‘Nuts and Bran’ which contains only nuts (N) and bran (B). The company has RM15000 to spend on nuts and bran. The cost of nuts is RM1200 and RM800 for bran.

     a.   Draw an isocost curve showing the different combinations of nuts and bran that the company can purchase.

(10 marks)

    b.   What will hapeen to the original isocost line if:

  1.     the capital reduce by 25%?
  2.     the price of nuts increase by RM200?
  3.     the price for both nuts and bran reduce by 15%?
1
Expert's answer
2022-01-17T16:10:13-0500

a.

cost:

"1200x+800y=15000"




b.

1.

"1200x+800y=15000(1-0.25)"

"1200x+800y=11250"


2.

"(1200+200)x+800y=15000"

"1400x+800y=15000"


3.

"1200(1-0.15)x+800(1-0.15)y=15000"

"1020x+680y=15000"